Sara and Melissa are best friends. Sara is 3 years older than Melissa. their combined age is 15. How old is Sara
Answer by ikleyn(52781)

.

Sara = x

Melissa = x-3


    x + (x-3) = 15

    2x - 3    = 15

    2x        = 15 + 3 = 18

     x                 = 18/2 = 9.


ANSWER.  Sara is 9 years old.

Solved.